OpenStudy (anonymous):

Find the solution of this equation. 1/3y=-6

jimthompson5910 (jim_thompson5910):

is it \[\Large \frac{1}{3}y = -6\] or \[\Large \frac{1}{3y} = -6\]

OpenStudy (anonymous):

The first 1

jimthompson5910 (jim_thompson5910):

ok how do you isolate y

jimthompson5910 (jim_thompson5910):

keep in mind that \[\Large \frac{1}{3}y = -6\] is the same as \[\Large \frac{y}{3} = -6\]

OpenStudy (anonymous):

Im not sure how to do it. that is why i asked.

jimthompson5910 (jim_thompson5910):

y/3 means "y divided by 3"

jimthompson5910 (jim_thompson5910):

how do we undo division by 3?

OpenStudy (anonymous):

Would it be 1 ?

jimthompson5910 (jim_thompson5910):

what's the opposite of division?

OpenStudy (anonymous):

Multiplication

jimthompson5910 (jim_thompson5910):

so you multiply both sides by 3 to undo division by 3

OpenStudy (anonymous):

So it would be 3 Wouldn't It ?

jimthompson5910 (jim_thompson5910):

y/3 = -6 multiply both sides by 3 to get y = -6*3 y = ??

OpenStudy (anonymous):

-18

jimthompson5910 (jim_thompson5910):

yep